# Mixed-Strategy Nash Equilibrium for Crowd Navigation

### v2

@@ Abstract
We address the problem of finding mixed-strategy Nash equilibrium for crowd navigation. Mixed-strategy Nash equilibrium provides a rigorous model for the robot to anticipate uncertain yet cooperative human behavior in crowds, but the computation cost is often too high for scalable and real-time decision-making. Here we prove that a simple iterative Bayesian updating scheme converges to the Nash equilibrium of a mixed-strategy social navigation game. Furthermore, we propose a data-driven framework to construct the game by initializing agent strategies as Gaussian processes learned from human datasets. Based on the proposed mixed-strategy Nash equilibrium model, we develop a sampling-based crowd navigation framework that can be integrated into existing navigation methods and runs in real-time on a laptop CPU. We evaluate our framework in both simulated environments and real-world human datasets in unstructured environments. Our framework consistently outperforms both non-learning and learning-based methods on both safety and navigation efficiency and reaches human-level crowd navigation performance on top of a meta-planner.
@@ I Introduction
Figure 1: Comparison of optimality criterion for decision-making in different navigation frameworks. (a) In traditional navigation, the robot makes optimal decisions, such as to minimize the risk of collision, in a given and non-interactive environment; (b) Cooperative navigation finds optimal cooperative decisions for both the robot and the human. For decisions made based on pure-strategy Nash equilibrium, the robot expects deterministic actions from humans, which is too assertive considering the uncertain nature of human behavior; (c) Our cooperative navigation framework uses mixed-strategy Nash equilibrium as the optimality criterion, which finds probabilities of actions that represent the optimal cooperation strategies between the robot and human. The probabilistic model of mixed-strategy maintains uncertainty during interaction.
@@
The ability to navigate fluently and safely in human-populated spaces is becoming increasingly crucial for deploying robot autonomy in real-world environments, examples include autonomous driving in populated areas [1], service robots at mass events [2], and industrial robots operating alongside human workers [3]. This problem of efficient robot navigation in human crowds with minimal disruption to humans is often referred to as crowd (social) navigation.
@@
Human behavior studies [4, 5] reveal that humans anticipate each other’s collision avoidance behavior when navigating in crowds. Empirical results further indicate that navigation algorithms that predict human cooperation during planning could improve both safety and navigation efficiency [6]. On the other hand, coupling the prediction of cooperative human behavior with motion planning is different from multi-agent planning, such as through crowd motion simulation [7, 8], where humans follow pre-defined action rules. Instead, both the humans and the robot should be modeled as optimal planners, whose individual objectives depend on others‘ actions [9]. As a result, the simultaneous prediction of cooperative human behavior and planning of robot motion become the outcome of assessing the coupled optimal actions of all the agents. This model leads to the application of game theory models in crowd navigation, specifically the application of Nash equilibrium [10] as an optimality criterion for decision-making.
@@
In a Nash equilibrium, no agent wants to change their strategy unless others do the same, since each agent’s strategy is optimal given other agents’ Nash equilibrium strategies. There are two kinds of strategies that can create a Nash equilibrium, pure-strategy and mixed-strategy, and we are interested in the latter. Pure-strategy refers to choosing deterministic and specific actions, while mixed-strategy refers to randomly sampling actions from a distribution of pure-strategies. Crucially, John Nash proved that not all games have a pure-strategy Nash equilibrium, but at least one mixed-strategy Nash equilibrium exists for all games [11]. Our interest in mixed-strategy Nash equilibrium stems from not only its mathematical rigor but also its non-deterministic nature. Prediction based on pure-strategy Nash equilibrium would expect humans to follow the paths exactly as predicted, which is too assertive given the uncertain nature of human behavior. Mixed-strategy Nash equilibrium, however, maintains such behavioral uncertainty during the robot’s decision-making. A high-level comparison of traditional navigation and cooperative navigation with pure-strategy Nash equilibrium and mixed-strategy Nash equilibrium is shown in Figure 1.
@@
Despite its potential, applying mixed-strategy Nash equilibrium to crowd navigation faces several challenges. First, computing mixed-strategy Nash equilibrium is often considered impractical due to its computation burden—even for a 3-agent game, the computation can be close to NP-complete [12]. Second, existing Nash equilibrium frameworks often focus on discrete actions and a finite number of strategies (e.g., poker). However, crowd navigation studies physical agents navigating in a continuous space, with extra physical constraints. Third, the practical implication of Nash equilibrium heavily depends on how well the game design matches human behavior, yet there is no consensus for a proper game model for crowd navigation.
In this work, we formally prove that—by modeling an agent’s mixed-strategy as a probabilistic belief of the agent’s intended path—a simple iterative Bayesian updating scheme guarantees the convergence to the global Nash equilibrium. In the corresponding game, each agent aims to minimize the expected risk of collision with other agents, as well as the deviation from a nominal mixed-strategy. Furthermore, we propose a framework to characterize the nominal mixed-strategy of each agent by extracting a temporal Gaussian process from offline trajectory datasets. This data-driven characterization process anecdotally captures the difference in crowd navigation patterns between humans and mice. Based on these theoretical results, we propose a sampling-based crowd navigation algorithm. The algorithm has a lower time complexity with respect to the number of agents compared to the state-of-the-art dynamic game solvers and runs in real-time on a laptop CPU. Given measurements of pedestrian positions and velocities, the algorithm infers mixed-strategy Nash equilibrium for up to 8 agents as a distribution of potential cooperation collision avoidance behaviors and chooses the optimal navigation decision for the robot accordingly. The algorithm is further integrated into a complete navigation stack for crowd navigation in unstructured real-world environments with static occupancy grid maps. Figure 2 provides a high-level illustration of the proposed iterative Bayesian updating scheme for computing mixed-strategy Nash equilibrium in crowd navigation.
@@
Our work diverges from existing game-theoretic crowd navigation methods. Existing methods follow a top-down approach, where no assumption about game objective structure and game decision space is made in advance—these components will be specified later for applications like crowd navigation. While the solution methodologies would apply to any game, not just crowd navigation game, this generality comes at the price of local optimal solvers [^1] Global Nash equilibrium solvers for arbitrary games is computationally intractable. and unaffordable computation cost and complexity for real-time decision-making in human crowds. In contrast, we take a bottom-up approach, where we formulate a specific game based on unique behavioral features in crowd navigation, such as goal-oriented and collision avoidance-driven behavior. This bottom-up approach provides a sufficient structure to develop a planner with game theory guarantees, including recovering global mixed-strategy Nash equilibrium, without compromising the computation efficiency for real-time decision-making. As demonstrated in our experiments, the computation efficiency of our method enables rapid replanning of the robot’s paths, which plays a critical role in improving navigation performance when interacting with potentially non-cooperative pedestrians.
@@
The rest of the paper is organized as follows: Section II reviews the literature on crowd navigation and discusses the key difference between our method and existing methods. We first introduce the iterative Bayesian updating scheme in Section III. Then we show how this iterative updating scheme is guaranteed to converge to a Nash equilibrium of a mixed-strategy crowd navigation game in Section IV, with extra safety property analysis of the Nash equilibrium. Section V introduces the method to learn nominal strategies from trajectory datasets. Section VI discusses how to use the mixed-strategy Nash equilibrium to infer human anticipation in practice with sampling-based approaches. Section VII contains the details of the evaluation results. Lastly, we conclude the paper and provide further discussion in Section VIII.
@@ II Related work
Figure 2: A high-level illustration of our framework. From left to right: The perception stack first estimates the states of the agents (e.g. robot and human positions). Then with the temporal Gaussian process kernel learned from offline datasets, we construct the initial mixed-strategies for the agents as Gaussian processes. After that, we apply the iterative Bayesian updating scheme to compute the mixed-strategy Nash equilibrium.
@@ II-B Prediction-then-planning methods
The limitations of conventional navigation methods in human crowds motivate researchers to develop human-aware navigation methods. One prevalent framework in this category is to predict human motion and plan robot actions to avoid hindering human motion, we name such framework as prediction-then-planning. Human motion can be predicted in the form of cost maps, such as from inverse reinforcement learning [21, 22] or unsupervised learning [23]. Human motion can also be directly predicted in forms of trajectories with deep learning methods, such as using long-short term memory models (LSTM) [24], generative adversarial networks (GAN) [25] or graph-based models [26]. We refer readers to [27] for a comprehensive review of human trajectory prediction. Note that, [28] reports that a simple constant velocity model could outperform deep learning-based trajectory prediction methods, arguing that existing neural network architectures are insufficient at capturing interpersonal interaction. In addition, another thread of research aims to develop fast reactive planners around dynamic obstacles, such as through field representations [29] or graph search [30].
Planning on top of human motion prediction also introduces the necessity of uncertainty-aware planning: motion planning needs to take into account predictive uncertainty in human motion. In [31], closed-loop belief updating of dynamic obstacles’ states is incorporated into receding horizon planning, which reduces the anticipated obstacle uncertainty and generates less conservative navigation behavior. A confidence-ware motion planning framework is proposed in [32] that maintains a Bayesian belief of the human motion prediction confidence with online observations. By explicitly modeling the confidence of prediction, the planner can be robust against unexpected human behavior. Robust motion planning with uncertain human motion prediction has also been investigated in other works. In particular, [33] introduces a decoupled framework that combines learning-based trajectory prediction [26] with risk-sensitive model predictive control [34]. Lastly, a flipped approach is introduced in [35] in the context of autonomous driving, where a risk-sensitive trajectory prediction approach is proposed for robust planning.
@@
While demonstrating more compliant navigation behavior alongside humans, decoupled prediction and planning is limited by its core assumption—that the robot’s action will not influence humans’ actions. Failing to account for human reaction could lead to robot actions that surprise humans, who in turn react out of the robot’s expectation, resulting in short oscillatory interaction, often referred to as “reciprocal dance” [36]. Failing to account for human reaction in uncertainty quantification could also lead to over-conservative navigation robot behavior—without incorporating human reaction to belief updating during planning, the over-approximated predictive uncertainty could lead the robot to consider all viable paths are unsafe and the only safe option is to stay still, a phenomenon often referred as the “freezing robot problem” [6].
@@ II-C Coupled prediction and planning
Given the limitations of decoupled prediction and planning, it becomes necessary to lift the assumption that the robot does not interfere with human intents. Thus, an alternative framework, named coupled prediction and planning, seeks to simultaneously plan robot actions and predict human actions. The seminal work [37] shows that cooperative collision avoidance (CCA)—where humans and robots collectively make decisions for collision avoidance—is crucial for avoiding artifacts such as “reciprocal dance” and “freezing robot problem”. The importance of CCA in dense crowds is experimentally verified in a psychophysical study [4].
@@
In [6], an individual’s intent is modeled as a Gaussian process and CCA is modeled as a joint decision-making process by coupling Gaussian processes through a collision avoidance-based likelihood function. The statistical optimality of coupled Gaussian processes is further investigated in [38, 39]. In [40], a distribution space crowd navigation model is proposed, with agent intent modeled as a distribution of trajectories. While similar to the mixed-strategy model in this work, the model in [40] is not a game-theory model but is instead a joint decision-making model similar to [6]. Topology-based abstractions are also used for modeling CCA, such as through braids [41, 42] or topological-invariance [43, 44, 45]. CCA is also studied from the perspective of opinion dynamics [46], with an explicit focus on breaking deadlock situations such as “reciprocal dance” [47].
@@
Another group of works focuses on the implicit modeling of cooperative collision avoidance. Implicit CCA models are often obtained from data from real-world or simulated human pedestrians—the availability of data will impact the choice of modeling technique. Given the cost of human data collection, implicit modeling with human data cannot afford directly training robot navigation policies with humans, instead uses techniques such as imitation learning [48] or inverse reinforcement learning [49, 50]. Instead of collecting human data, an alternative solution is to simulate cooperative navigation agents using decentralized multi-agent collision avoidance methods, such as [7, 8]. The benefit of implicit CCA modeling from simulated data is that the navigation policies can be trained directly with the reactive agents using reinforcement learning methods, such as [51, 52, 53, 54]. However, the sim-to-real transfer of such policies remains an open challenge, mainly due to the distribution shifts between simulated crowd behavior and real-world human behavior. It is also unclear whether the reinforcement learning policies explicitly predict human cooperation during planning, instead of treating humans merely as dynamic obstacles—the latter falls into the category of the prediction-then-planning framework.
@@
Note that there is a relevant group of works on social convention-aware navigation. The goal is to take into account social conventions during motion planning, such as people tend to walk in groups [55, 56] or people would prefer not to move when forming certain geometrical structures [57]. Though conceptually similar, we consider this body of work to be essentially different from coupled prediction and planning, because the social conventions are defined a priori. As a result, these planners do not necessarily take into account human reaction to the robot’s actions.
Limitation of existing coupled-prediction-and-planning methods
@@
The main limitation of existing coupled-prediction-and-planning methods is the lack of a rigorous cooperative model, given humans and the robot are cooperating through only mutual observation instead of explicit communication. Cooperation is often granted in coupled-prediction-and-planning methods. For example, methods that formulate crowd navigation as a joint decision-making problem assume that humans and robots share the same joint decision-making objective. We argue that each agent in crowd navigation should be modeled as an independent decision-making individual, the phenomenon of cooperative collision avoidance is an emergent behavior from each agent optimizing their individual objective, which partially depends on other agents’ actions. This insight naturally leads to the modeling of crowd navigation through game theory.

### v4

# Mixed Strategy Nash Equilibrium for Crowd Navigation

@@ Abstract
Robots navigating in crowded areas should negotiate free space with humans rather than fully controlling collision avoidance, as this can lead to freezing behavior. Game theory provides a framework for the robot to reason about potential cooperation from humans for collision avoidance during path planning. In particular, the mixed strategy Nash equilibrium captures the negotiation behavior under uncertainty, making it well suited for crowd navigation. However, computing the mixed strategy Nash equilibrium is often prohibitively expensive for real-time decision-making. In this paper, we propose an iterative Bayesian update scheme over probability distributions of trajectories. The algorithm simultaneously generates a stochastic plan for the robot and probabilistic predictions of other pedestrians’ paths. We prove that the proposed algorithm is equivalent to solving a mixed strategy game for crowd navigation, and the algorithm guarantees the recovery of the global Nash equilibrium of the game. We name our algorithm Bayes’ Rule Nash Equilibrium (BRNE) and develop a real-time model prediction crowd navigation framework. Since BRNE is not solving a general-purpose mixed strategy Nash equilibrium but a tailored formula specifically for crowd navigation, it can compute the solution in real-time on a low-power embedded computer. We evaluate BRNE in both simulated environments and real-world pedestrian datasets. BRNE consistently outperforms non-learning and learning-based methods regarding safety and navigation efficiency. It also reaches human-level crowd navigation performance in the pedestrian dataset benchmark. Lastly, we demonstrate the practicality of our algorithm with real humans on an untethered quadruped robot with fully onboard perception and computation.
@@ I Introduction
Figure 1: Comparison of optimality criterion in different navigation frameworks. (a) In traditional robot navigation, the robot makes optimal decisions, such as minimizing the risk of collision, in a given and non-interactive environment; (b) Cooperative navigation finds optimal cooperative decisions for both the robot and the human. With the pure strategy Nash equilibrium model, the robot expects deterministic actions from humans, which is too assertive considering the uncertain nature of human behavior; (c) Our cooperative navigation framework uses mixed strategy Nash equilibrium as the optimality criterion, which finds probabilities of actions that represent the optimal cooperation strategies between the robot and human. This model maintains uncertainty during the interaction.
@@
The ability to navigate fluently and safely in human-populated spaces is becoming increasingly crucial for deploying robots in real-world environments. Examples include autonomous driving in populated areas [1], service robots at mass events [2], and industrial robots operating alongside human workers [3]. This problem of safe and efficient robot navigation in human crowds with minimal disruption to humans is often referred to as crowd navigation.
@@
Human behavior studies [4, 5] reveal that humans anticipate each other’s collision avoidance behavior when navigating crowds. Empirical results further indicate that navigation algorithms that predict human cooperation for collision avoidance during planning could improve both safety and navigation efficiency [6]. On the other hand, coupling the prediction of cooperative human behavior with motion planning is different from multi-agent planning, such as through crowd motion simulation [7, 8], since humans do not follow pre-defined action rules. Instead, both the humans and the robot should be modeled as optimal planners whose individual objectives depend on others‘ actions [9]. As a result, the simultaneous prediction of cooperative human behavior and robot motion planning becomes the outcome of assessing the coupled optimal actions of all the agents. This design leads to the application of game theory models in crowd navigation, specifically Nash equilibrium [10] as an optimality criterion for the coupled prediction and planning process.
In game theory, each agent tries to optimize a game strategy—a plan of action—for their individual objective of the game. However, each agent’s individual objective depends on other agents’ strategies, so no agent can optimize their strategy in isolation. Nash equilibrium describes a balanced state among all agents’ strategies, where no agent wants to change their strategy unless others also change. In other words, each agent’s Nash equilibrium strategy is optimal given other agents’ Nash equilibrium strategies. Crowd navigation can be modeled as a game where the robot and pedestrians negotiate free space to ensure the joint safety of all agents and also minimize the compromise on their individual navigation plan. By finding the Nash equilibrium of such a game formula, the robot can plan a safe path while anticipating humans’ reactions, leveraging potential human cooperation to avoid over-cautious collision avoidance behaviors.
@@
There are two kinds of strategies that can lead to a Nash equilibrium: pure strategy and mixed strategy. We are interested in the latter. Pure strategy refers to choosing deterministic and specific actions, while mixed strategy refers to sampling actions from a probability distribution of possible pure strategies. Crucially, John Nash proved that not all games have a pure strategy Nash equilibrium, but at least one mixed strategy Nash equilibrium exists for all games [11]. Our interest in mixed strategy Nash equilibrium stems from not only its mathematical rigor but also its non-deterministic nature. Prediction based on pure strategy Nash equilibrium would expect humans to follow the paths exactly as predicted, which is too assertive given the uncertain nature of human behavior. Mixed strategy Nash equilibrium, however, maintains such behavioral uncertainty during the robot’s decision-making. A conceptual comparison of traditional navigation framework, cooperative navigation with pure strategy Nash equilibrium, and with mixed strategy Nash equilibrium is shown in Figure 1.
@@
Despite its potential, applying mixed strategy Nash equilibrium to crowd navigation faces several challenges. First, computing mixed strategy Nash equilibrium is often considered impractical due to its computation burden—even for a 3-agent game, the computation can be close to NP-complete [12]. Second, existing Nash equilibrium frameworks often focus on discrete actions and a finite number of strategies (e.g., poker). However, crowd navigation studies physical agents navigating in a continuous space. Third, the practicality of Nash equilibrium depends on how well the game design matches human behavior, yet there is no consensus on a game-theoretic model for crowd navigation.
In this work, we propose a simple iterative Bayesian update scheme for cooperative crowd navigation with mixed strategies. We formally prove that the algorithm guarantees the recovery of the global Nash equilibrium of a mixed strategy game suitable for crowd navigation. We name our algorithm Bayes’ Rule Nash Equilibrium (BRNE). In the BRNE game, each agent aims to minimize the expected risk of collision with other agents, while also minimizing the deviation of the optimal navigation strategy from a nominal navigation strategy. Furthermore, we propose a sampling-based model predictive crowd navigation framework based on BRNE, with the nominal strategies characterized based on Gaussian processes. The proposed navigation framework has a lower time complexity with respect to the number of agents compared to the state-of-the-art game-theoretic planners and runs in real-time on a laptop CPU and a low-power embedded computer. We further integrate the algorithm into an untethered quadruped robot and demonstrate the algorithm’s practicality with real humans using fully onboard perception and computation.
@@
Our work diverges from existing game-theoretic crowd navigation methods. Existing methods follow a top-down approach, where no specification of the game objective structure and the game decision space is made in advance—these components will be specified later for applications such as crowd navigation. While the solution methodologies would apply to any game, not just crowd navigation game, this generality comes at the price of local optimality [^1] Global Nash equilibrium solvers for arbitrary games is computationally intractable. and unaffordable computation cost for real-time decision-making in human crowds. In contrast, we take a bottom-up approach, formulating a specific game based on unique behavioral features in crowd navigation, such as goal-oriented and collision avoidance-driven behaviors. This bottom-up approach provides a sufficient structure to develop a planner with game theory guarantees, including recovering global mixed strategy Nash equilibrium, without compromising the computation efficiency for real-time decision-making, as demonstrated in our experiments.
@@
The rest of the paper is organized as follows: Section II reviews the literature on crowd navigation and discusses the key difference between our method and existing methods. We introduce the iterative Bayesian update scheme and the model predictive crowd navigation framework in Section III. Then we show how the algorithm guarantees the convergence to a global Nash equilibrium of a mixed strategy crowd navigation game in Section IV, with extra safety property analysis. Section V contains the details of the evaluation results and the real-world hardware demonstration. Lastly, we conclude the paper and provide further discussion in Section VI.
@@ II-B Prediction-then-planning methods
The limitations of conventional navigation methods in human crowds motivate researchers to develop human-aware navigation methods. One prevalent framework in this category is to predict human motion and plan robot actions to avoid hindering human motion, we name such framework as prediction-then-planning. Human motion can be predicted in the form of cost maps, such as from inverse reinforcement learning [21, 22] or unsupervised learning [23]. Human motion can also be directly predicted in forms of trajectories with deep learning methods, such as using long-short term memory models (LSTM) [24], generative adversarial networks (GAN) [25] or graph-based models [26]. We refer readers to [27] for a comprehensive review of human trajectory prediction. Note that, in [28], the authors report that a simple constant velocity model could outperform deep learning-based trajectory prediction methods, arguing that existing neural network architectures are insufficient to capture interpersonal interactions. In addition, another thread of research aims to develop fast reactive planners around dynamic obstacles, such as through field representations [29] or graph search [30].
Planning on top of human motion prediction also introduces the necessity of uncertainty-aware planning: motion planning methods that consider predictive uncertainty in human motion. In [31], a closed-loop belief update of dynamic obstacles’ states is incorporated into receding horizon planning, which reduces the anticipated obstacle uncertainty and generates less conservative navigation behavior. A confidence-ware motion planning framework is proposed in [32] that maintains a Bayesian belief of the human motion prediction confidence with online observations. The planner can be robust against unexpected human behavior by explicitly modeling prediction confidence. Other works have also investigated robust motion planning with uncertain human motion prediction. In particular, [33] introduces a decoupled framework that combines learning-based trajectory prediction [26] with risk-sensitive model predictive control [34]. Lastly, a flipped approach is introduced in [35] in the context of autonomous driving, where a risk-sensitive trajectory prediction approach is proposed for robust planning.
@@
While exhibiting more compliant navigation behavior alongside humans, decoupled prediction and planning is limited by its core assumption—that the robot’s action will not influence humans’ actions. Failing to account for human reaction could lead to robot actions that surprise humans, who in turn react out of the robot’s expectation, resulting in short oscillatory interaction, often referred to as “reciprocal dance” [36]. Failing to account for human reaction in uncertainty quantification could also lead to over-conservative navigation robot behavior—without incorporating human reaction to belief updating during planning, the predictive uncertainty could lead the robot to consider all viable paths are unsafe and the only safe option is to stay still, a phenomenon often referred to as the “freezing robot problem” [6].
@@ II-C Coupled prediction and planning
Given the limitations of decoupled prediction and planning, it becomes necessary to lift the assumption that the robot does not interfere with human intents. Thus, an alternative framework, named coupled prediction and planning, seeks to simultaneously plan robot actions and predict human actions. The seminal work [37] shows that cooperative collision avoidance (CCA)—where humans and robots collectively make decisions for collision avoidance—is crucial for avoiding artifacts such as the “reciprocal dance” and the “freezing robot problem”. The importance of CCA in dense crowds is experimentally verified in a behavioral study [4].
@@
In [6], an individual’s intent is modeled as a Gaussian process and CCA is modeled as a joint decision-making process by coupling Gaussian processes through a collision avoidance-based likelihood function. The statistical optimality of coupled Gaussian processes is further investigated in [38] and [39]. In [40], a distribution space crowd navigation model is proposed, with agent intent modeled as a distribution of trajectories. While similar to the mixed strategy model in this work, the model in [40] is not a game-theory model but is instead a joint decision-making model similar to [6]. Topology-based abstractions are also used for modeling CCA, such as through braids [41, 42] or topological-invariance [43, 44, 45]. CCA is also studied from the perspective of opinion dynamics [46], with a focus on breaking deadlock situations such as “reciprocal dance” [47].
@@
Another group of works focuses on the implicit modeling of cooperative collision avoidance. Implicit CCA models are often obtained from real-world data or simulated human pedestrians—the availability of data will impact the choice of modeling technique. Given the cost of human data collection, implicit modeling with human data cannot afford directly training robot navigation policies with humans; it uses techniques such as imitation learning [48] or inverse reinforcement learning [49, 50]. Aside from collecting human data, an alternative solution is to simulate cooperative navigation agents using decentralized multi-agent collision avoidance methods, such as [7] and [8]. The benefit of implicit CCA modeling from simulated data is that the navigation policies can be trained directly with the reactive agents using reinforcement learning methods, such as in [51], [52], [53], and [54]. However, the sim-to-real transfer of such policies remains an open challenge, mainly due to the distribution shifts between simulated crowd behavior and real-world human behavior. It is also unclear whether the reinforcement learning policies explicitly predict human cooperation during planning, instead of treating humans merely as dynamic obstacles—the latter case falls into the category of the prediction-then-planning framework.
@@
Note that there is a relevant group of works on social convention-aware navigation. The goal is to take into account social conventions during motion planning, such as people’s tendency to walk in groups [55, 56] or people’s preference for not moving when forming certain geometrical structures [57]. Though conceptually similar, we consider these works to be essentially different from coupled prediction and planning, because the social conventions are defined a priori. As a result, these planners do not necessarily consider humans’ reactions to the robot’s actions.
Limitations of coupled prediction and planning methods
@@
The main limitation of existing coupled prediction and planning methods is the lack of a rigorous cooperation model since humans and robots cooperate only through mutual observations instead of explicit communication. Cooperation is often granted in coupled prediction and planning methods. For example, methods that formulate crowd navigation as a joint decision-making problem assume that humans and robots share the same joint decision-making objective, making these methods more similar to a centralized multi-agent planning framework. We argue that each agent in crowd navigation should be modeled as an independent decision-making individual. Cooperative collision avoidance is an emergent phenomenon generated by agents’ desire to optimize their individual objectives, which depends on other agents’ actions. This insight naturally leads to the application of game theory to crowd navigation.